# Cooperative Cruising: Reinforcement Learning based Time-Headway Control for Increased Traffic Efficiency

### v1

@@ 4.3 Practical Design Choices
An AI system that influences CAVs equipped with technologies such as ACC must address several key objectives to be practical:
@@
  • Deployability and User Acceptance: For successful deployment, the system’s decisions must be transparent, instilling confidence among drivers and stakeholders.
To address safety, our system influences existing safety-certified ACC systems by sending them desired time-headway commands that can only increase the time-headway above its default value. We assume the availability of low-bandwidth vehicle-to-infrastructure communication that allows vehicles to receive a few floating-point values every few seconds representing desired time-headways, supporting the simplicity and generality objectives. The infrastructure is assumed to measure traffic metrics across different road segments, such as average speed, density, and throughput. This type of sensing is already available, and was recently used in a large-scale open-road experiment (Lee et al. 2024), supporting the deployability of our proposed system. The system’s time-headway decisions can be made transparent to users and support user-acceptance.
@@ 4.4 Fixed-Valued Time-Headway Control
Before proposing an RL-based time-headway control policy, it is natural to ask whether a simpler fixed-value time-headway control policy can outperform human-driven traffic, and if so, whether RL can offer significant additional improvements. We therefore design a controller that sends an optimized fixed time-headway signal to road segments located before a bottleneck. The controller activates when vehicles are detected on the merging road, and deactivates when no vehicles are present there. When deactivated, vehicles return to their default time-headway. The optimal time-headway value is optimized by a parameter sweep.
4.5 RL-based Safe Time-Headway Control
The dynamic nature of traffic suggests that a controller capable of adjusting time-varying headway values would outperform a fixed time-headway approach (Yanakiev and Kanellakopoulos 1995). The ability of reinforcement learning (RL) to learn from interactions with the environment makes it a good fit for complex, dynamic, and stochastic environments such as traffic control. We design an RL-based controller that continuously monitors the traffic state, determines time-headway commands for each controlled road segment, and communicates these commands to automated vehicles traveling within those segments. While RL policies typically do not have safety guarantees, our RL controller ensures safety by sending headway commands above the minimum safe threshold to a safety-certified commercial ACC system, which consistently maintains a safe following distance. Notably, no crashes occurred in our experiments.
To apply RL to our traffic control problem, we model the problem as a discrete-time, finite-horizon Markov Decision Process (MDP) (Puterman 2014), defined by a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},P,R,\rho_{0},T)$, where $\mathcal{S}$ is the set of possible environment states, $\mathcal{A}$ is the set of all possible actions, $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ is a state transition probability distribution, $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is a reward function mapping a given state and the action taken from it to a numeric reward, $\rho_{0}:\mathcal{S}\rightarrow[0,1]$ is a distribution over initial states, and $T$ is the problem’s time horizon. In an MDP, the goal of the RL algorithm is to learn a decision-making policy $\pi:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ that stochastically maps states to actions and maximizes the expected cumulative sum of rewards over all trajectories $E_{\tau}\sum_{t=0}^{T}r(s_{t},a_{t})$. Here $\tau$ is a trajectory $\left[s_{0},a_{0},s_{1},a_{1},...,s_{T},a_{T}\right]$, where the initial state $s_{0}$ is sampled from the initial state distribution: $S_{0}\sim\rho_{0}$, actions are sampled from the policy: $a_{t}\sim\pi\left(s_{t}\right)$, and the next state in the trajectory is sampled from the transition probability: $s_{t+1}\sim P\left(s_{t},a_{t}\right)$, defined by the traffic simulator.
@@ 4.5 RL-based Safe Time-Headway Control
To allow effective control of traffic density, our state representation encapsulates relevant information that is (i) needed for predicting the traffic’s evolution over time, and (ii) is reasonable to obtain using current technology. Therefore, a state includes the average speeds and densities in each of 21 road segments preceding and following the traffic bottleneck area, where each segment is about 100 meter long. In general, both speed and density are necessary in the state representation to avoid ambiguity.
@@
The actions $\mathcal{A}$ correspond to the required time-headway values for automated vehicles in each segment. Since density typically accumulates in the segments preceding the traffic bottleneck, changing density in these segments offers the highest potential for improving traffic efficiency through density control. Therefore, our action is a vector of real numbers representing desired time-headways in each controlled segment preceding the bottleneck. In our experiments we tried setups with 2-5 controlled segments. For the tested scenario, controlling the time-headway for 2 road segments was found to allow faster RL convergence without performance compromise. Thus, the RL algorithm results in this work were trained using 2 control segments. While more granular actions spaces could be used, such as separate headway for each lane in each controlled segment, a single headway per segment would be more practical to implement in the real-world, as it avoids the need to estimate vehicles’ lane-position.
@@
The reward function plays a crucial role in the RL training process, as it guides the agent toward maximizing the desired performance metric. As described in Section 4.1, our performance metric is the relative average speed increase compared with a baseline simulated human-driven traffic. However, the average speed of a vehicle can only be computed once a vehicle had completed its route, so using it would delay the reward and pose challenges for current RL algorithms.
To provide a more immediate reward that serves as a proxy for our metric, we use a time-delay reward function that is measured relative to free-flow traffic conditions. The reward at time $t$ is:
$r_{t}=\frac{1}{C}\sum_{i=1}^{N_{t}}\left(\frac{v_{t}^{\left(i\right)}-v_{free}\left(x_{t}^{\left(i\right)}\right)}{v_{free}\left(x_{t}^{\left(i\right)}\right)}\right)\Delta t$ (3)
where $N_{t}$ is the number of vehicles planned to enter the simulation by time $t$, $x_{t}^{\left(i\right)}$ and $v_{t}^{\left(i\right)}$ are the location and velocity of vehicle $i$ at time $t$, $v_{free}\left(x\right)$ is the speed limit of the road at location $x$, and $C$ is a normalization factor that scales episode returns to the interval $[0,1]$, to avoid numerical issues when using neural networks. For delayed vehicles which did not yet enter the simulation, the speed is assumed to be 0. When this reward is accumulated for all simulation time-steps, it provides an approximation for the average time delay for all vehicle trajectories, relative to free flow. This time-delay corresponds to our average speed performance metric as described in Section 4.1. The advantage of this reward over our metric is its immediate feedback on the impact of the current traffic state over the overall average speed.
@@ 5 Empirical Analysis
In this section, we describe the experimental setup used to evaluate the proposed time-headway control strategies and present the results obtained from the simulations.
Figure 3: Time-headway control performance: Simulation performance results for (a) single-lane scenarios, and (b) multi-lane scenarios. The performance of our safe RL-based controller and fixed-valued time-headway control baseline is measured relative to simulated human-driven traffic (dashed black line). Error bars show 95% confidence intervals for mean performance values, each derived from 30 simulations.
Experiments were conducted in the SUMO simulation environment using a highway network with a single merge, based on an approximately 2 km segment of route I-24 in Tennessee, USA (see Section 3.2). The route was divided into 20 segments of approximately 100 meters each, plus an additional segment for the merging road. This segment length was chosen to ensure effective sensing and control resolution while being long enough to maintain stable control policies. The default time-headway parameter value is 1.5 seconds, and the speed limit on the highway is 31.29 m/s (70 mph). Traffic inflow to the highway matches the maximum capacity of 1800 vehicles/hr/lane. Vehicles from the merging road enter the simulation at maximum inflow after a 200-second warm-up for a duration of 30 seconds in the single-lane scenarios, and 50 seconds in the multi-lane scenarios, to create realistic traffic disturbances that negatively affect traffic flow. The simulations ran for 500 seconds to allow congestion to form and dissipate. An action interval of 2.5 seconds was chosen to reflect a realistic duration for influencing traffic dynamics, considering vehicle response times and acceleration characteristics.
@@
    (Baseline) Mixed Traffic with Fixed-Valued Time-Headway Control: Traffic consisting of both human drivers and a fraction of 10-100% CAVs, which set their headway based on commands sent by the fixed-valued time-headway controller described in Section 4.4. The controller was tuned to the best performing fixed-valued time-headway.
@@
    (Ours) Mixed Traffic with RL-based Controller: Traffic consisting of both human drivers and a fraction of 20-100% CAVs, which set their headway based on commands sent by the RL-based time-headway controller described in Section 4.5.
@@
Figure 3 presents results from hundreds of experiments under the most challenging traffic flows and varying percentages of controlled vehicles. Each scenario configuration was run 30 times with different random seeds, enabling the computation of 95% confidence intervals. We focused on two scenario types: a simplified merge with a single-lane main road, and a complex merge with a multi-lane highway:
@@
Highway vehicles drive on the rightmost lane, no lane changes are allowed, and a single-lane road merges into the highway. This simplified setup tests the feasibility of time-headway control. The tuned fixed-value control improves merge efficiency and traffic flow over the human baseline across different CAV fractions (Figure 3a). RL-based control further enhances performance for CAV fractions of 40% and above, achieving up to a 16% average speed increase in the 100% CAV scenario.
@@
All highway lanes are used, lane-changes are allowed, and a single-lane road merges into the highway, creating a congestion-prone bottleneck. Time-headway control signals are sent to all lanes in controlled segments. While lane-specific control could be more effective, it requires extra sensing, limiting practicality. Fixed-value time-headway control improves merge efficiency and traffic flow only at low or 100% CAV fractions (Figure  3b). In contrast, our RL-based controller outperforms both the human baseline and fixed-value control across all CAV fractions, achieving up to a 7% increase in average speed at 100% CAV fraction.
Overall, our RL-based controller outperforms the fixed-headway controller in scenarios with complex dynamics: specifically, in single-lane settings with CAV fractions of 40% or more, and in multi-lane settings with CAV fractions of 20% or more, where fixed-headway controllers are challenging to tune and may struggle to represent effective control strategies. These experiments robustly demonstrate the effectiveness of our approach, which is the first to handle realistic multilane scenarios.
@@ 6 Conclusions
This paper proposes a dynamic time-headway control approach for increasing the average travel speed of vehicles while maintaining safety in high-volume highway traffic. By integrating with existing traffic estimation technology and low-bandwidth vehicle-to-infrastructure connectivity, and leveraging safety-certified adaptive cruise control systems, our method offers a practical path towards real-world implementation. Our safe reinforcement learning-based time-headway controller outperforms both baselines and alternative approaches across a variety of automated vehicle penetration rates, in both single- and multi-lane realistic, simulated scenarios featuring hundreds of vehicles. Notably, even at low penetration rates, adjusting time-headways led to measurable improvements in average traffic speeds.
@@ Appendix C Hyperparameters
$\frac{1}{C}$ (reward normalization) 1e-5

### v2

@@ 4.3 Practical Design Choices
An AI system for CAVs equipped with technologies such as ACC must address several key objectives to be practical:
@@
  • Deployability and User Acceptance: For successful deployment, the system’s decisions must be transparent to instill confidence in drivers and stakeholders.
To address safety, our system is designed to influence existing safety-certified ACC systems by sending them desired time-headway commands that can only increase the time-headway above its default value. To promote generality, we model a representative scenario of a typical highway merge. We simplify implementation by designing state and action spaces independent of vehicle count, and assume the availability of low-bandwidth vehicle-to-infrastructure communication, where vehicles periodically receive a few floating-point values representing desired time-headways. The infrastructure is assumed to measure traffic metrics across different road segments, such as average speed, density, and throughput. This type of sensing is already available, and was recently used in a large-scale open-road experiment (Lee et al. 2024), supporting the deployability of our proposed system. The system’s time-headway decisions can be made transparent to users and support user-acceptance.
@@ 4.4 Fixed-Valued Time-Headway Control
Before proposing an RL-based time-headway control policy, it is natural to ask whether a simpler fixed-value time-headway policy can outperform human-driven traffic and, if so, whether RL can provide significant additional improvements. We therefore design a controller that sends an optimized fixed time-headway signal to road segments located before a bottleneck. The controller activates when vehicles are detected on the merging road within 200 meters before the merge and deactivates when no vehicles are present in this segment. When deactivated, vehicles return to their default time-headway. We find the optimal time-headway value using a parameter sweep.
4.5 RL-based Time-Headway Control
The dynamic nature of traffic suggests that a controller capable of adjusting time-varying headway values would outperform a fixed time-headway approach (Yanakiev and Kanellakopoulos 1995). RL’s ability to learn from environment interactions makes it a good fit for complex, dynamic, and stochastic setups such as traffic control. We design an RL-based controller that continuously monitors the traffic state, determines time-headway commands for each controlled road segment, and communicates these commands to automated vehicles traveling within those segments. Although RL policies typically lack safety guarantees, our RL controller ensures safety by issuing headway commands above the minimum safe threshold to a safety-certified commercial ACC system, which consistently maintains a safe following distance. Notably, no crashes occurred in our experiments.
MDP
To apply RL to our traffic control problem, we model the problem as a discrete-time, finite-horizon Markov Decision Process (MDP) (Puterman 2014), defined by a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},P,R,\rho_{0},T)$, where $\mathcal{S}$ is the set of possible environment states, $\mathcal{A}$ is the set of all possible actions, $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ is a state transition probability distribution, $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is a reward function mapping a given state and the action taken from it to a numeric reward, $\rho_{0}:\mathcal{S}\rightarrow[0,1]$ is a distribution over initial states, and $T$ is the problem’s time horizon. In an MDP, the goal of the RL algorithm is to learn a decision-making policy $\pi:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ that stochastically maps states to actions and maximizes the expected cumulative sum of rewards over all trajectories $E_{\tau}\sum_{t=0}^{T}r(s_{t},a_{t})$. Here $\tau$ is a trajectory $\left[s_{0},a_{0},s_{1},a_{1},...,s_{T},a_{T}\right]$, where the initial state $s_{0}$ is sampled from the initial state distribution: $S_{0}\sim\rho_{0}$, actions are sampled from the policy: $a_{t}\sim\pi\left(s_{t}\right)$, and the next state in the trajectory is sampled from the transition probability $s_{t+1}\sim P\left(s_{t},a_{t}\right)$, defined by the traffic simulator.
@@ 4.5 RL-based Time-Headway Control
To enable effective traffic density control, our state representation encapsulates relevant information that (i) is necessary for predicting traffic dynamics over time and (ii) can be feasibly obtained with current technology. Specifically, the state includes average speeds and densities across 21 road segments before and after the traffic bottleneck, with each segment spanning approximately 100 meters. In general, both speed and density are essential in the state representation to prevent ambiguity.
@@
The actions correspond to the required time-headway values for automated vehicles in each segment. Since density tends to accumulate in segments preceding the traffic bottleneck, adjusting density in these segments offers the greatest potential for improving traffic efficiency through density control. Thus, our action is a vector of real numbers representing desired time-headways in each controlled segment before the bottleneck. Our experiments tested setups with 2–5 controlled segments, with 2 segments enabling faster RL convergence without performance loss. Consequently, the reported empirical results use this setup. While it is possible to explore more granular action spaces such as separate headways for each lane within a segment, they are less practical for real-world implementation as they require vehicle lane-position estimation.
@@
The reward function plays a crucial role in the RL training process, as it guides the agent toward maximizing the desired performance metric. As described in Section 4.1, our performance metric is the relative increase in average speed compared to a baseline of simulated human-driven traffic. However, the average speed of a vehicle can only be computed once a vehicle had completed its route, so using it would result in delayed reward which poses challenges for current RL algorithms.
To provide a more immediate reward, we use a time-delay reward function that is measured relative to free-flow traffic conditions. This reward is closely correlated to the performance metric when the traveled distance is fixed (see supplementary material), as is the case for the vast majority of vehicles in the simulation. The reward at time $t$ is:
$r_{t}=\frac{1}{C}\sum_{i=1}^{N_{t}}\left(\frac{v_{t}^{\left(i\right)}-v_{free}\left(x_{t}^{\left(i\right)}\right)}{v_{free}\left(x_{t}^{\left(i\right)}\right)}\right)\Delta t,$ (3)
where $N_{t}$ is the number of vehicles planned to enter the simulation by time $t$, $x_{t}^{\left(i\right)}$ and $v_{t}^{\left(i\right)}$ are the location and velocity of vehicle $i$ at time $t$, $v_{free}\left(x\right)$ is the speed limit of the road at location $x$, and $C$ is a normalization factor that scales episode returns to the interval $[0,1]$, to avoid numerical issues when using neural networks. For delayed vehicles which did not yet enter the simulation, the speed is assumed to be 0. When this reward is accumulated for all simulation time-steps, it provides an approximation for the average time delay for all vehicle trajectories, relative to free flow. The advantage of this reward is its immediate feedback on the impact of the current traffic state over the overall average speed.
@@ 5 Empirical Analysis
In this section, we describe the experimental setup for evaluating the proposed time-headway control strategies and present results from hundreds of large-scale simulations.
Experiments were conducted in the SUMO simulation environment on a realistic 2 km, four-lane highway with a merging road (described in Section 3.2). The highway was divided into 20 segments of approximately 100 meters each, plus an additional segment for the merging road. This segment length balances effective sensing and control resolution while ensuring stable control policies.
The default time-headway is 1.5 seconds, and the highway speed limit is 31.29 m/s (70 mph). Traffic inflow matches the highway’s maximum capacity of 1800 vehicles/hr/lane. Vehicles from the merging road enter at maximum inflow after a 200-second warm-up, allowing the simulation to reach a steady state. Merging traffic continues for 30 seconds in single-lane scenarios and 50 seconds in multi-lane scenarios, creating realistic traffic disturbances that impact flow.
Simulations ran for 500 seconds to allow congestion to develop and dissipate. An action interval of 2.5 seconds was chosen to reflect a realistic time frame for influencing traffic dynamics, considering vehicle response times and acceleration characteristics.
@@
    (Baseline) Mixed Traffic with Fixed-Value Time-Headway Control: Traffic consists of both human-driven vehicles and 20-100% CAVs, which follow headway commands from the fixed-value time-headway controller (described in Section 4.4). The controller was tuned to use the best-performing fixed time-headway.
@@
    (Ours) Mixed Traffic with RL-based Controller: Traffic consists of both human-driven vehicles and 20-100% CAVs, which follow headway commands from the RL-based time-headway controller (described in Section 4.5).
@@
Figure 3 presents results from hundreds of experiments under varying percentages of controlled vehicles. We used maximal vehicle inflow, which is the most sensitive to traffic disturbances. Each scenario configuration was run 30 times with different random seeds, enabling the computation of 95% confidence intervals. We focused on two scenario types: a simplified merge into a single-lane road and a complex merge into a multi-lane highway.
@@
Highway vehicles travel on a single-lane road with another single-lane road merging into it. This simplified setup tests time-headway control feasibility. The tuned fixed-value control improves merge efficiency and traffic flow over the human baseline across different CAV fractions (Figure 3a). RL-based control further enhances performance for 40% CAV fractions and above, achieving up to a 13% average speed increase at 100% CAV fraction.
@@
Highway vehicles travel on a four-lane road, with a single-lane road merging into it, creating a congestion-prone bottleneck. Time-headway control signals are sent to all lanes in controlled segments. While lane-specific control could be more effective, it requires extra sensing, limiting practicality. Fixed-value time-headway control improves merge efficiency and traffic flow only at low or 100% CAV fractions (Figure  3b). In contrast, our RL-based controller outperforms both the human baseline and fixed-value control across all CAV fractions, achieving up to a 7% increase in average speed at 100% CAV fraction.
Overall, our RL-based controller outperforms the fixed-headway controller in scenarios with complex dynamics: specifically, in single-lane settings with CAV fractions of 40% or more, and in multi-lane settings with CAV fractions of 20% or more, where fixed-value headway controllers struggle to represent effective control strategies. These experiments demonstrate the effectiveness of our approach, which is the first to handle realistic multilane scenarios. Notably, the RL controller is most effective at high CAV fractions. With the widespread adoption of technologies such as ACC, it could become a viable tool for mitigating highway congestion.
Figure 3: Time-headway control performance: Simulation performance for (a) single-lane, and (b) multi-lane scenarios. Performance of our safe RL-based controller and fixed-valued time-headway control baseline is measured relative to simulated human-driven traffic (dashed black line). Error bars show 95% confidence intervals for mean performance values, each derived from 30 simulations.
@@ 6 Conclusions
This paper proposes a dynamic time-headway control approach for increasing the average travel speed of vehicles while maintaining safety in high-volume highway traffic. By integrating with existing traffic estimation technology and low-bandwidth vehicle-to-infrastructure connectivity, and leveraging safety-certified adaptive cruise control systems, our method offers a practical path towards real-world implementation. Our safe reinforcement learning-based time-headway controller outperforms both baselines and alternative approaches across a variety of automated vehicle penetration rates, in both single- and multi-lane realistic simulated scenarios featuring hundreds of vehicles. Notably, even at low penetration rates, adjusting time-headways led to measurable improvements in average traffic speeds.
@@ Appendix C Hyperparameters
$\frac{1}{C}$ (reward normalization) $10^{-5}$
@@
Appendix D Reward Function Analysis
The average velocity of vehicle $i$ is
$\bar{v}_{i}=\frac{L_{i}}{T_{i}},$ (4)
where $T_{i}$ and $L_{i}$ are the total travel time and the total travel distance of vehicle $i$, respectively. Thus, assuming the travel distance (route) of each of the vehicles is fixed, increasing the average speed amounts to minimizing $T_{i}$.
For a vehicle $i$ in the simulation, the total travel time can be computed by:
$T_{i}=\sum_{t=1}^{T_{sim}}a_{i}\left(t\right)\text{d}t,$ (5)
where $T_{sim}$ is the total number of simulation timesteps, $\text{d}t$ is the duration of each timestep, and $a_{i}\left(t\right)$ is a boolean parameter which is 1 if vehicle $i$ was planned to enter the simulation before time $t$ and it did not yet exit the simulation before time $t$, and 0 otherwise. This boolean parameter is crucial, since without it the controller can just prevent vehicles from entering the simulation, and thus decrease the road density and increase average velocity for a smaller number of vehicles without penalty. Note, that since we measure $L_{i}$ only on the travel distance within the simulated road, we assume here that the velocity of delayed vehicles is 0. While this is not accurate, this still provides a better approximation for the average speed of the full road than disregarding delayed vehicles completely. This will also provide higher penalties on delaying vehicles from entering the simulation.
Therefore, the function to minimize is the average total travel time of vehicles in the simulation:
$\bar{T}=\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{T_{sim}}a_{i}\left(t\right)\text{d}t,$ (6)
where $N$ is the number of vehicles that are expected to enter the simulation.
Reinforcement learning algorithms train a controller to maximize the expected value of the return, which is the time-aggregation of the (discounted) reward:
$\max_{\pi}\lim_{T_{sim}\rightarrow\infty}E\left\{\sum_{t=1}^{T_{sim}}\gamma^{t}r\left(t\right)\right\},$ (7)
where $\pi$ is the policy, and $\gamma<1$ is the discount factor. For this theoretical analysis, we assume $\gamma=1$, and note that by replacing the return in Equation 7 with the (negative) average travel time in Equation 6, we derive the following reward function:
$r(t)=-\frac{1}{N}\sum_{i=1}^{N}a_{i}\left(t\right)\text{d}t.$ (8)
However, using the total travel time reward has a few drawbacks:
  1. 1.
    Delayed reward – The reward for an action is delayed since it only depends on the number of vehicles exiting the simulation. Therefore, only when a vehicle leaves the simulation, the penalty is reduced. In this case, distinguishing between actions that decrease travel time and actions that increase it is a challenge, since the total travel time depends on the entire time series of actions for each vehicle.
  2. 2.
    Start / end point dependent – Positive reward is achieved only for vehicles that exited the simulation. This means that if we increase the number of simulated road segments, the reward will be further delayed.
To deal with that, we seggest a time-delay reward function:
$r(t)=-\frac{1}{N}\sum_{i=1}^{N}a_{i}\left(t\right)\left(1-\frac{v_{i}\left(t\right)}{v_{free}\left(x_{i}\left(t\right)\right)}\right)\text{d}t,$ (9)
where $v_{i}\left(t\right)$ and $x_{i}\left(t\right)$ are the velocity and position of vehicle $i$ at time $t$ respectively, and $v_{free}$ is the free-flow velocity at a certain location (i.e., the speed limit there). This is an immediate reward that uses real-time velocity readings of the vehicles. Therefore, actions get immediate feedback based on the magnitude of the travel delay they caused.
The sum of these rewards over the entire simulation duration gives the average time delay measured from the free flow completion time:
$\sum_{t=1}^{T}r\left(t\right)=$ $-\frac{1}{N}\sum_{t=1}^{T_{sim}}\sum_{i=1}^{N}a_{i}\left(t\right)\left(1-\frac{v_{i}\left(t\right)}{v_{free}\left(x_{i}\left(t\right)\right)}\right)\text{d}t$
$=$ $-\frac{1}{N}\sum_{t=1}^{T_{sim}}\sum_{i=1}^{N}a_{i}\left(t\right)\left(1-\frac{v_{i}\left(t\right)\text{d}t}{v_{free}\left(x_{i}\left(t\right)\right)}\right)$
$=$ $-\frac{1}{N}\sum_{t=1}^{T_{sim}}\sum_{i=1}^{N}a_{i}\left(t\right)\left(1-\frac{\text{d}x_{i}\left(t\right)}{v_{free}\left(x_{i}\left(t\right)\right)}\right)$
$=$ $-\frac{1}{N}\sum_{t=1}^{T_{sim}}\sum_{i=1}^{N}a_{i}\left(t\right)\left(1-\text{d}t_{free}^{\left(i\right)}\left(t\right)\right)$
$=$ $-\frac{1}{N}\sum_{i=1}^{N}\left(T_{i}-T_{free}^{\left(i\right)}\right)$
$=$ $-\frac{1}{N}\sum_{i=1}^{N}\Delta T_{i},$ (10)
where $\text{d}x_{i}\left(t\right)$ is the distance vehicle $i$ traveled during timestep $t$, $T_{free}^{\left(i\right)}$ is the travel time of vehicle $i$ assuming it drove at free-flow speed, and $\Delta T_{i}$ is its time delay. $\text{d}t_{free}^{\left(i\right)}\left(t\right)$ is the time it would take vehicle $i$ to travel a distance of $\text{d}x_{i}\left(t\right)$ in free-flow. The last steps in the derivation above assume vehicle $i$ completed its entire journey. Since each vehicle’s travel distance is its route length
$\sum_{t=1}^{T_{sim}}a_{i}\left(t\right)\text{d}x_{i}\left(t\right)=L_{i},$ (11)
the free flow completion time of the entire route is
$T_{free}^{\left(i\right)}=\sum_{t=1}^{T_{sim}}a_{i}\left(t\right)\frac{\text{d}x_{i}\left(t\right)}{v_{free}\left(x_{i}\left(t\right)\right)}=\sum_{t=1}^{T_{sim}}a_{i}\left(t\right)\text{d}t_{free}^{\left(i\right)}(t)$ (12)
which is constant for all vehicles with the same route. This computation further assumes that free flow velocity is constant between $x_{i}\left(t\right)$ and $x_{i}\left(t\right)+\text{d}x_{i}\left(t\right)$ for all values of $t$.
It is important to note that as long as vehicle $i$ completes its entire route, maximizing the expected value of the above objective leads to maximizing the expected value of the (weighted) relative change in average velocity compared to a baseline simulation:
$E\left\{-\frac{1}{N}\sum_{i=1}^{N}\left(T_{i}-T_{free}^{\left(i\right)}\right)\right\}$
$=$ $E\left\{-\frac{1}{N}\sum_{i=1}^{N}\left(T_{i}-T_{base}^{\left(i\right)}+T_{base}^{\left(i\right)}-T_{free}^{\left(i\right)}\right)\right\}$
[…]
$\propto$ $E\left\{-\frac{1}{N}\sum_{i=1}^{N}\left(T_{i}-T_{base}^{\left(i\right)}\right)\right\}$
$=$ $E\left\{-\frac{1}{N}\sum_{i=1}^{N}T_{i}\left(1-\frac{T_{base}^{\left(i\right)}}{T_{i}}\right)\right\}$
$=$ $E\left\{-\frac{1}{N}\sum_{i=1}^{N}T_{i}\left(1-\frac{\bar{v}_{i}}{\bar{v}_{base}^{\left(i\right)}}\right)\right\}$
$=$ $E\left\{\frac{1}{N}\sum_{i=1}^{N}T_{i}\left(\frac{\bar{v}_{i}-\bar{v}_{base}^{\left(i\right)}}{\bar{v}_{base}^{\left(i\right)}}\right)\right\},$ (13)
where $\bar{v}_{base}^{\left(i\right)}$ and $T_{base}^{\left(i\right)}$ are the average velocity of vehicle $i$, and its total travel time in the baseline (human-only) simulation, respectively. This is similar to the metric used in the paper, up to a weight factor for each vehicle which is equal to the total travel time for that vehicle. This means that the reward function puts more weight on vehicles whose total travel time is large. However, since the total travel time of the vehicles depends on the policy, it cannot be easily normalized in the reward function.